OpenStudy (anonymous):

(-1,1) from rectangular to a polar coordinates. a. (2, 315degrees) b. (sqrt2, 135degrees) c. (sqrt2, 315degrees) d. (sqrt2, 225degrees)

OpenStudy (anonymous):

\[\Large r=\sqrt{x^2+y^2}\] and \[\Large tan(\theta)=\frac{y}{x}\] use x=-1 and y=1 to get r and theta.

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

@sami-21 I'm not getting it :/

OpenStudy (anonymous):

\[\Large r=\sqrt{(-)^2+(1)^1}\] \[\Large r=\sqrt{2}\] \[\Large tan(\theta)=\frac{1}{-1}\] \[\Large \theta=tan^{-1}(-1)\]

OpenStudy (anonymous):

-45

OpenStudy (anonymous):

Am I right? So that means...

OpenStudy (anonymous):

it is in the second quadrant !!!! (-1,1) so 180+(-45)=135

OpenStudy (anonymous):

Oh o.k. I understand. Thank you so much, I really appreciate your help